OpenStudy (anonymous):

the population is 300,000 with a 3% annual growth rate how long until the population will reach 1.2 million?

OpenStudy (tkhunny):

?? \(300000(1.03)^{t} = 1200000\) Use your best logarithms to solve for 't'. It's a WHOLE LOT less than 100 years.

OpenStudy (anonymous):

After 1 year, population = 300,000 x 1.03 after 2 years, population = 300,000 x 1.03 x 1.03 = 300,000 x (1.03)^2 after 3 years, population = 300,000 x 1.03 x 1.03 x 1.03 = 300,000 x (1.03)^3 Note the pattern forming. After n years, population = 300,000 x (1.03)^n (^ is to the power of) Problem now written as: 300,000 x (1.03)^n = 1,200,000 divide both sides of equation by 300,000 1.03^n = 4 take ln of both sides (ln button calculator) ln(1.03^n) = ln(4) now use logarithm rule lna(b^n) = n x lna(b) This rule can be proven for example with log10(10^2) = 2 x log10(10). Log10 is log button on calc. So n x ln (1.03) = ln(4) divide both sides by ln (1.03) n = ln(4) divided by ln(1.03) calculator does this. Answer = 46.8995445 years. So after 47 years of annual growth rate of 3%, population reaches 1.2 million. A check of this answer is to sub solution into the original problem. 300,000 x (1.03)^46.8995445 = 1.2 million Note about using logarithms to solve. I always find it hard to remember what way to write these problems so I type log10 (1000) in the calculator to work out the form in which to write the problem. Note log10(1000) = 3. This means 10^3 = 1000 l